### v1

# Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents

Abstract
Uncertainty quantification (UQ) for large language models (LLMs) is a key building block for safety guardrails of daily LLM applications. Yet, even as LLM agents are increasingly deployed in highly complex tasks, most UQ research still centers on single-turn question-answering. We argue that UQ research must shift to realistic settings with interactive agents, and that a new principled framework for agent UQ is needed. This paper presents the first general formulation of agent UQ that subsumes broad classes of existing UQ setups. Under this formulation, we show that prior works implicitly treat LLM UQ as an uncertainty accumulation process, a viewpoint that breaks down for interactive agents in an open world. In contrast, we propose a novel perspective, a conditional uncertainty reduction process, that explicitly models reducible uncertainty over an agent’s trajectory by highlighting “interactivity” of actions. From this perspective, we outline a conceptual framework to provide actionable guidance for designing UQ in LLM agent setups. Finally, we conclude with practical implications of the agent UQ in frontier LLM development and domain-specific applications, as well as open remaining problems.
[…]
1 Introduction
Figure 1: Comparison between UQ setups. Traditional LLM UQ (a) measures the uncertainty of an answer given a question, whereas the UQ for LLM reasoning (b) expands the problem by considering multi-step responses rather than a single response. Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau$-bench Airline; barres2025tau).
LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent] at the next level of autonomy [Steinberger2026OpenClaw]. In such settings, failures are no longer limited to incorrect text generation: agents may act prematurely under unresolved ambiguity, propagate errors across long interaction trajectories, or commit to outcomes that are costly or difficult to undo. To be deployed responsibly, agents must be able to assess and act upon the likelihood of failure [kochenderfer2015decision]. This makes reliable uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
Despite the importance, most existing UQ research treats LLMs as static oracles: a system examined in isolation, prompted once, and evaluated by the uncertainty of a single response or a chain of responses (Figure 1, left). These methods implicitly assume a static system in which no new information is acquired after the initial prompt [malinin2021uncertainty, kadavath2022language, farquhar2024detecting]. Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation. While appropriate for single-turn inference and non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
Interactive agents operate in multi-turn interactive environments with feedback from users, tools, and external systems (Figure 1, right). For example, a flight-booking agent may need to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences. Early uncertainty should prompt information-seeking actions, while subsequent dialogue and database queries can resolve ambiguities and reduce uncertainty, enabling confident execution. Here, uncertainty does not strictly accumulate–it may also decrease as a direct consequence of interaction, even as the trajectory grows longer.
In this paper, we argue that making LLM agents trustworthy requires a shift in focus: from pointwise uncertainty of answers to structured uncertainty flow in open, interactive decision processes; and a new framework is necessary for this. We present the following perspective:
Uncertainty quantification for an LLM agent must incorporate interactivity to model reducible uncertainty: in non-isolated, agentic systems, total uncertainty should be allowed to decrease as new information is acquired through interaction.
To ground this perspective in a formal discussion, we first provide a concrete definition of agent UQ and a general formulation that captures broad categories of existing UQ setups (Sec. 3). Specifically, we model the agent’s problem-solving trajectory as a stochastic process over actions, observations, and environment states, represented as a simple graphical model. Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ approaches arise as special cases.
This unified view also exposes a fundamental limitation of existing methods: uncertainty is implicitly treated as accumulating monotonically along the trajectory, regardless of the role played by different actions; thus fails to relate uncertainty to a task reward. To address this limitation, we introduce in Sec. 4 a conceptual framework that models agent UQ as a conditional uncertainty reduction process, through an information-gating mechanism that operationalizes the distinction between uncertainty-increasing and uncertainty-reducing actions by considering their interactivity. We further derive the analytic bounds that improve the interpretability of the resulting uncertainty estimates.
In addition, Sec. 5 discusses practical implications of agent UQ across diverse domains, including healthcare, software engineering, and robotics, as well as frontier LLM research on adaptive reasoning and post-training. Finally, we outline open problems unique to uncertainty modeling in agentic systems (Sec. 6). Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware agentic systems.
  • We formalize agent UQ as a distinct problem setup and show why existing LLM UQ formulations are insufficient for interactive agent setups.
  • To tackle that limitation, we present a conceptual framework for agent UQ by viewing it as a conditional uncertainty reduction process.
  • To give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
2 Related Work
UQ in LLMs.
In contrast to classic UQ [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs. Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over samples [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal]. While promising, they mostly use single-step question-answering as a testbed, which limits their practical applicability to agentic setups.
Figure 2: Graphical model for an agent problem-solving trajectory with examples. Given a task specification $E_{0}$ and an initial user query $O_{0}$, an agent spans a multi-turn trajectory characterized by a chain of action AA, observation $O$, and environment state $E$. This simple abstraction describes some representative agentic prompting methods such as ReAct [yao2022react]. See Appendix A.1 for details.
UQ in reasoning and agent setups.
Beyond single-step generation, researchers show growing interest in measuring the confidence of LLMs during their reasoning. This kind of (un)certainty information has been used to improve the quality of response [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning]. Recently, there have also been attempts to model uncertainty in interactive and/or agentic setups [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal]. Although UQ here aims to model uncertainty dynamics over long-horizon generation, existing methods do not capture the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new framework for agent UQ.
3 A General Formulation for Agent UQ
3.1 Problem Statement and Definitions
Imagine a flight booking agent depicted in Figure 1 right. Here, the types of action $A$ and observation $O$ are diverse: for $A$, there is information-seeking, questioning for clarification, calling an API function, and thinking or updating without interactive behaviors. The observation $O$, on the other hand, can be a database status report, tool-calling results, or the user’s clarification. Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., code.
Concretely, given a task specification $E_{0}$, e.g., a flight booking, a user makes a stochastic initial query $O_{0}$, and the agent first requests the user’s ID to get relevant information. Then, it calls a programmed function to get privileged information from the database. Not only does the agent interact with the user or tool, but it also thinks about better future actions (i.e., planning) or makes a system update without external observation $O=\{\varnothing\}$. All interactions and other actions are stored in environment state $E$ for context-grounded actions over a long horizon inference. This makes up a trajectory $\mathcal{F}_{\leq t}$ (chain of $A$, $E$, and $O$) up to $t$-turn which continues until the task-specific termination rule is met, e.g., when the agent submits an answer [yang2018hotpotqa, jimenez2024swebench, wei2025browsecomp], the user generates a stop sign [yao2024tau, barres2025tau], or goal completion [shridhar2021alfworld] with multiple constraints [xie2024travelplanner].
Now, to formalize UQ in this real-world agent context, we begin with a concrete definition of the agent system in Definition 1, and introduce a graphical model for the agent trajectory in Figure 2. This can generalize many existing agent benchmark setups [mialon2023gaia, yao2024tau, yang2018hotpotqa, song2023llm, jimenez2024swebench, xie2024travelplanner, liu2024agentbench, wei2025browsecomp].
Definition 1 (Stochastic Agent System).
Let $E_{i}$ be an environment state, a mixture of a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and a state of the system database, which is not directly accessible to the agent (only partially observable via tool-calling). Let $O_{i}$ and $A_{i}$ be the $i$-th turn observation and action derived from distributions $P$ and $P_{\pi,\mathcal{T}}$, respectively, where $\mathcal{T}$ indicates a tool set and $\pi$ is an LLM policy. Then, under a task specification $E_{0}$ and the user’s initial query $O_{0}$, we define the trajectory as a sequence of turns $\mathcal{F}_{\leq T}=\{(A_{t},E_{t},O_{t})\}_{t=0}^{T}$ where $\mathcal{F}_{0}=(E_{0},O_{0})$. Given a deterministic update function $h(\cdot)$, the generative process for each step $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
To express this trajectory with a simple mathematical representation, we introduce a dynamic Bayesian network [murphy2002dynamic] in Figure 2. With this graphical model, dependency structures of $A$, $E$, and $O$ can be explained as follows: (1) since the environment state $E_{i-1}$ contains the entire conversation history of $A_{\leq i-1}$ and $O_{<i-1}$ up to the i1i-1 step [^1] An adaptive memory structure [xu2025mem, xiong2025memory] can replace this ever-growing memory., the current action $A_{i}$ depends only on $E_{i-1}$ and $O_{i-1}$; (2) the current observation $O_{i}$ from the environment also depends only on $A_{i}$ and $E_{i}$; (3) the transition of the environment state ($E_{i-1}\rightarrow E_{i}$) can be defined as a deterministic function [^2] In general, this system transition can be also stochastic, but we consider a simplified yet still realistic scenario. $h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system state given the current action.
Under this Bayesian network, the joint probability for the trajectory can be factorized as below,
$P(\mathcal{F}_{\leq T})$ $=P(E_{0},O_{0})\Pi_{i=1}^{T}P(\mathcal{F}_{i}|\mathcal{F}_{i-1})$
$=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$
where $\mathcal{F}_{0}=(E_{0},O_{0})$, and $P(E_{i}|\cdot)=1$ as it is determistic. Now, we define the UQ problem for this stochastic agent system as follows.
Definition 2 (Agent UQ).
Let $U(X)\geq 0$ be an uncertainty function that takes a random variable $X$ or its distribution to produce a non-negative real value. Given a $T$-turn trajectory, agent UQ aims to estimate both a turn-level uncertainty $U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ for t=1,,Tt=1,...,T and a trajectory-level uncertainty $U(\mathcal{F}_{\leq T})=U(\mathcal{F}_{0},...,\mathcal{F}_{T})$, as a joint total uncertainty.
In the machine learning field, the most popular choices of the uncertainty function $U(X)$ are Shannon entropy [shannon1948mathematical] $\mathbb{E}[-\log P(X)]$, negative log probability $\log P(X=x)$, and their variants. These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suit for sequential modeling. For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix A and B for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
$U(\mathcal{F}_{\leq T})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$
$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$
where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ denotes an initial query uncertainty plus a task volatility.
Desideratum.
While previous LLM UQ setups usually focus on conditional uncertainty of a final answer, agent UQ aims at estimating joint uncertainty over the full trajectory, where the estimate from a calibrated agent should be predictive of the reward on that trajectory. Formally, for all $r_{1}>r_{2}$,
$\mathbb{E}\big[U(\mathcal{F}_{\leq T})|r(\mathcal{F}_{\leq T})=r_{1}\big]<\mathbb{E}\big[U(\mathcal{F}_{\leq T})|r(\mathcal{F}_{\leq T})=r_{2}\big],$ (1)
where $r(\cdot)$ denotes a real-value reward function commonly set as a success-failure binary annotator.
We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus. See Appendix C for discussion.
3.2 A Unified View on Existing UQ Setups
We now show that our agent UQ formulation captures various existing UQs as special cases.
Reduction to single-step LLM UQ.
When $t=1$ and the action space is restricted to responses, LLM UQ can be cast as a special case
$U(\mathcal{F}_{\leq T})=U(O_{0})+U(A_{1}|O_{0})\geq U(A_{1}|O_{0}),$
where $O_{0}$ and $A_{1}$ denote the given question and the corresponding model answer, respectively. Most current LLM UQ literature addresses only the conditional answer uncertainty $U(A_{1}|O_{0})$ [malinin2021uncertainty, fadeeva2023lm, farquhar2024detecting]. A few also consider the question ambiguity $U(O_{0})$ [hou24decomposing, tomov2025the] to reflect the reality of wild user query [min2020ambigqa].
Reduction to multi-step reasoning UQ.
When the action space involves multi-step reasoning (e.g., chain-of-thought, wei2022chain), we have
$U(\mathcal{F}_{\leq T})$ $=U(O_{0})+\sum_{i=1}^{T}U(A_{i}|A_{<i},O_{0})$ (2)
$\geq\sum_{i=1}^{T}U(A_{i}|A_{<i},O_{0})$ (3)
$\geq\max_{i\in\{1,...,T\}}U(A_{i}|A_{<i},O_{0})$ (4)
$\geq\sum_{i=1}^{T}w_{i}U(A_{i}|A_{<i},O_{0}),$ (5)
where $w_{i}\geq 0$ and $\sum_{i=1}^{T}w_{i}=1$. Here, $O_{0}$ and $A_{T}$ denote the given initial query and the model’s final response, and $A_{i},...,A_{T-1}$ are intermediate outputs. Similar to the single-step setup, most existing multi-step UQs [tanneru2024quantifying, tao2025revisiting] do not consider the question ambiguity $U(O_{0})$, so that reduces Eq. 2 to Eq. 3, while a few consider it [leang2025picsar]. Meanwhile, rather than summing up all the step-wise uncertainties, some approaches consider the most uncertain one, i.e., the lowest confidence used in fu2025deep, which can be represented by Eq. 4. Finally, Eq. 5 expresses various weighted average methods, such as a standard length-normalized uncertainty [kang2025scalable] by setting $w_{i}=\frac{1}{T}$, and a tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies, i.e., top-K mean [fu2025deep] or model-based estimates [zhao2025uncertainty].
Connection to process reward modeling.
We further draw a connection between this trajectory-level UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward. For example, lightman2024lets proposed the product of the step-wise probabilities and the minimum probability between steps as the aggregation method. These are analogs to Eq. 3 and Eq. 4, respectively.
Figure 3: Limitation of existing UQ (a) and our suggestion (b). Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure. We urge expanding a dimension of interest, interactivity, and moving on to reducible uncertainty modeling of agents.
Limitation of existing frameworks.
From this unified view, we see that they all model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn. One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa. However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward. See Fig. 3 (a), a scenario wherein a user does not fully specify their goal upfront. After searching for options, the agent either reasons itself or interacts with the user to delegate decisions. Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence. Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 1).
4 How to Model Uncertainty Dynamics of Interactive Agents?
4.1 Unique Challenge
In contrast to the LLM UQ, agent UQ spans a multi-turn trajectory in an open environment. This presents a unique challenge: modeling cascade uncertainty in an interactive system. Here, the desired UQ method should be able to precisely aggregate uncertainties across the entire trajectory and account for additional information gained through interactions. While there is emerging research on step-wise uncertainty aggregation [kang2025scalable, fu2025deep], they examine an isolated, noninteractive setting where an agent does not receive any external feedback during inference (e.g., Fig. 1, b). Meanwhile, agent UQ considers a setup full of interactions with a user and an environment, making it a non-isolated, interactive system.
Since an agent in this open world can gather additional information, make clarifications, and request confirmations at each step by interacting with the user and external databases through conversations and tool calls, the uncertainty dynamics in this interactive system includes not only propagations of turn-level uncertainty, but also reductions. As shown in Fig. 3, the existing perspective does not consider “interactivity” in agent UQ and prohibits reducible uncertainty modeling. In this paper, we argue that we must introduce interactivity in agent UQ to model bi-directional uncertainty flow.
4.2 Agent UQ as a Conditional Uncertainty Reduction Process
We need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, depending on the agent’s interactivity at each step. While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec. 3.
Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
$U(\mathcal{F}_{\leq{T}})=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})]$
$=U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})[1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}]$
$\geq U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})g(\mathcal{F}_{i}).$ (6)
Here, we introduce a conditional gating function $g(\cdot)$ that reduces or increases a current turn uncertainty depending on the action. That is, for a set of valid uncertainty reduction actions, $\mathcal{A}^{-}$, we define:
$g(\mathcal{F}_{i})=\begin{cases}\frac{-\text{Info}(O_{i};O_{0}|E_{i}\backslash O_{0})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{if}\penalty 10000\ A_{i}\in\mathcal{A}^{-},\\ 1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{otherwise.}\end{cases}$
where $\text{Info}(\cdot;\cdot)$ denotes mutual information or pointwise mutual information that measures the amount of information gain by having the current observation $O_{i}$ w.r.t. initial query $O_{0}$ given conversation history $E_{i}\backslash O_{0}$. For each turn, if the agent’s action is classified as an interactive (inviting a user or a tool) and evidential one (factual or not in conflict with stored data), then $g(\cdot)$ produces the amount of information gain with negative sign to realize uncertainty reduction; otherwise, it propagates uncertainty (See Fig. 4).
Figure 4: Illustration on the proposed agent UQ paradigm. We propose the conditional uncertainty reduction process for LLM agents by discerning interactive and evidential actions from others.
Importantly, the signed gating function handles uncertainty in a more interpretable manner than traditional estimates by letting not only quantitative but also directional interpretations. Besides, this uncertainty model allows us to derive the analytic extrema, i.e., maximum and minimum, of total uncertainty in closed forms as shown in Lemma 1.
Action Category Interactivity Example
Information-gathering interactive Agent uses a read tool (e.g., get_reservation_details or search_flights) to retrieve data or asks the user for missing information, e.g., “Could you provide your reservation number and last name?”.
Asking clarification or confirmation to user interactive Agent asks the user preferences or confirmation on decision, such as, “Do you want me to proceed with booking FQ8APE?”.
Thinking non-interactive Agent plans the future action sequence based on the previous trajectory.
State-changing tool call (writing) non-interactive Agent calls the tools that modify the database (e.g., cancel_reservation, book_reservation, update_reservation_flights) commit to an outcome.
Providing final information to user non-interactive Agent reports the result of an action, such as “Your reservation has been cancelled; your refund will be processed” or “Your flight has been rebooked to SFO departing at 8 a.m.”.
Table 1: Example action classes in Airline booking scenario in $\tau$-bench. A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty. See Appendix E for extended analyses.
Lemma 1 (Extrema of Information Gating).
Let the lower bound of agent total uncertainty in Eq. 6 be $\tilde{U}(\mathcal{F}_{\leq T})$, denote $U(X):=H(X)=\mathbb{E}[-\log P(X)]$ and $\text{Info}(X;Y):=I(X;Y)=\mathbb{E}[\log\frac{P(X,Y)}{P(X)P(Y)}]$, then, we have:
$\tilde{U}(\mathcal{F}_{\leq T})$ $\geq H(E_{0},O_{0})-\sum_{i=1}^{T-1}I(O_{i},O_{0}|E_{i}\backslash O_{0})+H(A_{T}|E_{T-1},O_{T-1}),$
$\tilde{U}(\mathcal{F}_{\leq T})$ $\leq H(E_{0},O_{0})+\sum_{i=1}^{T}H(A_{i},O_{i}|E_{i-1},O_{i-1}).$
Proof.
If $A_{i}\in\mathcal{A}^{-}$ (resp. $A_{i}\notin\mathcal{A}^{-}$) for all $1\leq i\leq T-1$, then $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction (resp. propagation) process, deriving above inequalities (See Appendix F for details). ∎
Interpretation.
Lemma 1 states that an agent’s total uncertainty can be expressed as an ever-decreasing process starting from a task-specific initial uncertainty reduced by the amount of extra information delivered by current observation, given the previous trajectory till its final behavior in the ideal case; whereas it is continually stacked up by the amount of joint entropy between action and observation per turn in worst case. Together with the directional interpretation turn-level uncertainty (increase or decrease), these analytic extrema help us better understand the numerical uncertainty, which is usually hard to interpret [van2019communicating, belem2024perceptions, kirchhof2025self].
Sketch of implementation.
To operationalize this idea, we first need an action classifier to determine whether $A_{i}\in\mathcal{A}^{-}$ or not. One can adopt a model-based evaluator [kim2025beyond] and a rule-based verifier [zeng2025reinforcing] to assess the interactivity and evidentiality of actions, respectively, where $A_{i}\in\mathcal{A}^{-}$ if and only if $A_{i}$ is both evidential and interactive. An example categorization of actions in an airline assistant task is provided in Table 1. After that, four different quantities need to be estimated: (1) initial query uncertainty, (2) action uncertainty, (3) observation uncertainty, and (4) mutual information between the query and observation given the previous trajectory. For these estimations, we can leverage the output probability [malinin2021uncertainty, aichberger2024rethinking], consistency in multiple trials [farquhar2024detecting], world modeling [chae2025web, anonymous2026dreamphase], and a learning neural estimator [molavipour2020conditional, gritsai2025mist]. We leave the actual implementation for future work. See Appendix D for additional notes.
5 Practical Implications
Developing an agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments. We outline implications for frontier LLM research and three specialized domains: healthcare, coding, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
5.1 Advancing Frontier LLMs
Test-time scaling through reasoning is now a de facto standard technique to enhance the quality of LLM responses on challenging tasks [wei2022chain, muennighoff2025s1]. However, there is still huge room for improvement, e.g., mitigating overthinking and/or pursuing efficiency [sui2025stop, aggarwal2025optimalthinkingbench]. Agent UQ provides concrete guidance to develop an adaptive reasoning method, i.e., early stopping or interaction invoking, via uncertainty budgeting. In addition, multi-turn reinforcement learning (RL) has emerged as a promising yet challenging research theme in eliciting the agentic behavior and reasoning of LLMs [zhou2024archer, shani2024multi, zeng2025reinforcing]. Agent UQ helps tackle the main bottlenecks, credit assignment and when-to-explore problems [thrun1992efficient, pecka2014safe, choudhury2025process], limited positive feedback [lee2025banel, lee2026semi], and inspires an information pursuit policy [chattopadhyay2022interpretable] by leveraging directional and quantitative uncertainty information as feedback.
5.2 Reliable Clinical Decision Supporting Agents
Modern LLM agents offer a promising vision for healthcare, moving beyond static knowledge retrieval to autonomous systems with complex reasoning and tool-execution capabilities [wang2025survey]. However, the transition from promising prototypes to reliable clinical agents faces critical obstacles stemming from the system’s inability to recognize and manage its own limitations. We believe that agent UQ can fill this gap. For example, although ferber2025development observed a huge potential of tool-calling LLM agents in oncology diagnoses, the remaining error rate underscores the necessity of a “human-in-the-loop” workflow. Here, agent UQ can serve as a gatekeeper, e.g., flagging highly accumulated uncertainty moments to invite humans, automating during low-uncertainty periods, while pursuing total uncertainty reduction for the right final diagnosis. This aligns with the concept: adaptive healthcare that envisions a paradigm shift from passive prediction to active, risk-aware inquiry [hinostroza2025ai].
5.3 Reliable Software Engineering Agents
Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for an agent UQ framework. Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, or commit to a fix. Besides, considering uncertainty-triggered rollback and branching algorithm, in which an agent treats increasing uncertainty after an action as a signal to revert or branch by mirroring the checkpoint-rollback workflows of human engineers. These highlight the potential of UQ for reliable agentic coders.
5.4 Relible Embodied Agents in Cyber-Physical Systems
Embodied agents in cyber-physical systems face heterogeneous uncertainty from sensing, dynamics, and human intent [li2024embodied, fung2025embodied, team2025gemini], and they actively reduce uncertainty through interaction by re-sensing the environment or interacting with humans before committing to irreversible physical actions [xu2025compliant]. For example, a robot instructed to retrieve a fragile object may explore or seek clarification rather than execute a grasp under high uncertainty. This makes embodied autonomy a canonical instance of a conditional uncertainty reduction process, where information-gathering actions reduce uncertainty and state-changing actions increase commitment and risk. By explicitly quantifying and managing this uncertainty dynamic, our agent UQ solution offers safer action selection [zhang2025safevla], principled deferral [hagenow2025realm], and robust execution [romer2025failure].
6 Open Problems
Although we have so far presented a conceptual roadmap and the practical implications of agent UQ, significant open problems remain as follows.
Developing rigid benchmarks at scale.
As in the case of process reward modeling mentioned in Section 3.2, annotating across long-horizon trajectories introduces non-trivial efforts and costs [choudhury2025process]. Moreover, as true uncertainty is unmeasurable, i.e., only a proxy-based evaluation is possible, establishing benchmarks for agent UQ stands on the next level of difficulty than the static LLM evaluation [ielanskyi2025addressing]. Investigation on defining proper agent performance measures [kapoor2025holistic, hendrycks2025definition] with scalable evaluation and annotation methods for agent rollouts [yang2024react, chang2024agentboard, kim2025beyond] would be a crucial stepping stone.
Long-horizon black-box uncertainty estimation.
LLM agents are mostly offered by proprietary services, where we cannot access the output probability of the agent. While model owners and providers can leverage low-cost white-box UQ approaches, others may have to rely on expensive black-box UQ that requires heavy sampling per action over a long trajectory [zhang2024luq]. This poses a great challenge for low-resource users in developing UQ with state-of-the-art agents.
Solution multiplicity.
Is high uncertainty per step due to the agent’s lack of knowledge about the right action for that step? Or, is it due to the intrinsic multiplicity of valid actions in that step? We cannot identify which of them is the right source of high uncertainty! Besides, in contrast to the classic question answering, the task bid to an agent is too complex to be perfectly specified at once. Thus, each intermediate step is intrinsically ambiguous and allows multiple valid actions, making it impossible to specify the source of the high uncertainty.
UQ in multi-agent or self-evolving agent systems.
While we confine our setup to single-agent, non-evolving agent systems, there has been remarkable advancement in multi-agent systems [guo2024large, wu2024autogen] and self-evolving agents [wang2024voyager, jiang2025adaptation]. Modeling the joint uncertainty dynamics of multiple agents or considering the non-stationary toolkit and agent parameters over time brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
7 Conclusion
We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the bi-directional dynamics model of uncertainty. We laid the foundation for this by providing a concrete definition and general formulation of agent UQ that abstracts a wide variety of existing UQ approaches. To tackle the limitations of existing framing, we presented a novel perspective, conditional uncertainty reduction process, with actionable guidelines and a theoretical analysis. As a concluding remark, we quote Schrödinger’s note on living in physics [schrodinger1974what]: “Every living means an increase in entropy of the part of the world, and it can only survive by continually drawing negative entropy from its environment.”
Limitations
Although this paper provided a novel high-level vision for agent UQ with a crystallized formulation, we built this vision on a relatively simple setup with some assumptions: reliable observation $O_{i}$ from the environment and user, and deterministic environment update $E_{i}\rightarrow E_{i+1}$. However, agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment, where our current proposal does not ensure its effectiveness. In that scenario, we may need to explicitly model uncertainty in the agent’s evidence stream as well, e.g., trust/ignorance about tool outputs, retrieval, and user-provided facts, by leveraging formalisms, such as subjective logic [jsang2018subjective, cheng2020there], imprecise probability theory [walley1991statistical], and formal belief representation [formalbelief22], to safely manage memory and act under unreliable observations in stochastic world.
Ethical Considerations
This work aims to establish a foundation for uncertainty quantification of LLM agents, which can directly contribute to society, science, and human well-being [han2023recipe, ren2025towards, oh2025flex, hendrycks2025definition]. We believe that the realization of our proposed framework can significantly boost the reliability of the LLM agent in a dynamic, interactive inference setup, thereby facilitating risk mitigation in high-stakes decision-making. However, we also acknowledge that the proposed framework can be misused by allowing adversaries to efficiently steer the agent towards its blind spots or manipulate uncertainty thresholds to bypass safety filters [wang2025agentvigil]. Therefore, we recommend alleviations such as limiting exposure to attack-enabling signals, red-teaming, and monitoring for abuse, which avoid overpromising what low uncertainty implies.
Acknowledgments
We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, and Kyungwoo Song for their sharp feedback on the draft. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23- 1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Paul Bogdan acknowledges the support by the National Science Foundation (NSF) under the NSF Award 2243104 under the Center for Complex Particle Systems (COMPASS), NSF Mid-Career Advancement Award BCS-2527046, U.S. Army Research Office (ARO) under Grant No. W911NF-23-1-0111, Defense Advanced Research Projects Agency (DARPA) Young Faculty Award and DARPA Director Fellowship Award, Okawa foundation award, National Institute of Health (NIH) under R01 AG 079957, and Intel faculty awards.
Appendix
We cover the following content in this Appendix.
  • (A) Additional Details on Formulation
  • (B) Alternative Uncertainty Measures
  • (C) Extended Context
  • (D) Note on Implementation
  • (E) Action Categorization
  • (F) Missing Proof
  • (G) Discussion and Future Work
Appendix A Additional Details on Formulation
A.1 Expression of various agentic prompting methods under our formulation.
In Section 3.1, we have introduced a definition of stochastic agent system (Def. 1) with a probabilistic graphical model of specific conditional dependency assumptions 2. Under this model, we can abstract some representative agentic prompting methods [ahn2022can, yao2022react, shinn2023reflexion, patil2025berkeley] with a unified language yet distinctive conditions as below.
  1. 1.
    Vanilla function calling [patil2025berkeley]: Same as Definition 1. Stochastic process of $(A_{t},E_{t},O_{t})$ without any extra conditions.
  2. 2.
    Few-shot prompting [ahn2022can]: Just augmenting initial task-specification variable as $E^{{}^{\prime}}_{0}=E_{0}\cup\{\mathcal{F}^{(i)}\}_{i=1}^{N}$ with some example rollouts $\mathcal{F}^{(i)}=\{(A_{t},E_{t},O_{t})\}_{t=1}^{T^{i}}$, which may implicitly affect the action transition probability, $P_{\pi,\mathcal{T}}(\cdot)$, by conditioning.
  3. 3.
    ReAct [yao2022react, wang2024executable]: Adding a constraint on the action sampling as $A_{i}\sim P_{\pi,\mathcal{T}}(A^{\texttt{thk}}|E_{i-1},O_{i-1})$ if $A_{i-1}\in\mathcal{A}^{\neg\texttt{thk}}$ else $A_{i}\sim P_{\pi,\mathcal{T}}(A^{\neg\texttt{thk}}|E_{i-1},O_{i-1})$, where $A^{\texttt{thk}}\in\mathcal{A}^{\texttt{thk}}$ and $A^{\neg\texttt{thk}}\in\mathcal{A}^{\neg\texttt{thk}}$ randon varaibles from the partitions of $\mathcal{A}$ standing for the space of textual reasoning and all others, respectively.
  4. 4.
    Reflexion [shinn2023reflexion]: Do multiple trials $\mathcal{F}^{(\leq i)}=\{\mathcal{F}^{(1)},...,\mathcal{F}^{(i)}\}$ until $r^{(i)}:=r(\mathcal{F}^{(i)})=1$ or reach maximum trials, otherwise augmenting the per-trial initial state with an external long-term memory $E^{(i+1)}_{0}=E^{(i)}_{0}\cup\{\mathcal{M}(\mathcal{F}^{(j)},r^{(j)})\}_{j\leq i}$, where the long-term memory $\mathcal{M}(\cdot)$ constructs with textual feedback generated from an LLM prompted by the current trial trajectory and a corresponding binary reward. Thus, the action transition probability, $P_{\pi,\mathcal{T}}(\cdot)$, is continually evolved by reflecting feedback from historical runs.
Overall, this suggests that our agent formulation is general and flexible enough to capture a broad class of representative agentic prompting scenarios.
A.2 Uncertainty Instantiations and Total Uncertainty Expansion.
After defining the agent UQ in Def 2, we presented a trajectory-level total uncertainty that expresses the joint uncertainty across multiple turns in an additive form of the initial query uncertainty, action uncertainty, and observation uncertainty.
Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
  1. 1.
    Information content (negative log probability), $U(X=x):=-\log P(X=x)$, to measure a point-wise surprisal for a given observation.
  2. 2.
    Entropy, $U(X):=H(X)=\mathbb{E}[-\log P(X)]$, to measure the expected amount of surprisal.
  3. 3.
    Relative entropy, $U(X):=\mathbb{D}_{\rm KL}\big(Q(X)||P(X)\big)$ with a pre-defined reference distribution $Q(X)$ and Kullback-Leibler divergence $\mathbb{D}_{\rm KL}$, to measure the expected amount of surprisal given a prior knowledge (if we have any).
For information content, it is trivial to show by just taking the negative logarithm to $P(\mathcal{F}_{\leq T})$. For the entropy and relative entropy, we have the following chain rule [cover1999elements],
  • $H(\mathcal{F}_{0},...,\mathcal{F}_{T})=\sum_{i=0}^{T}H(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})$,
  • $\mathbb{D}_{\rm KL}\big(Q(\mathcal{F}_{0},...,\mathcal{F}_{T})||P(\mathcal{F}_{0},...,\mathcal{F}_{T})\big)=\sum_{i=0}^{T}\mathbb{D}_{\rm KL}\big(Q(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})||P(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})\big)$,
which directly drives our uncertainty expansion for $U(\mathcal{F}_{\leq T})$ in Section 3.1.
Appendix B Alternative Uncertainty Measures
Although we have mainly focused on the above three information-theoretic uncertainty measures throughout the paper, there are plenty of alternatives one can consider depending on the problem setups. In this section, we examine some possible alternative uncertainty measures, Rényi entropy, Tsallis entropy, and informational energy.
Rényi entropy [renyi1961measures]
has been widely applied in modern machine learning, ecology, biodiversity science, and quantum information, among other fields. Given an order parameter $0<\alpha<\infty$ and $\alpha\neq 1$, it is defined as $H_{\alpha}(X)=\frac{1}{1-\alpha}\log(\sum_{x\in\mathcal{X}}P(x)^{\alpha})$ and generalize many other entropies: max-entropy $H_{0}(X)=\log|\mathcal{X}|$, Shannon entropy $\lim_{\alpha\rightarrow 1}H_{\alpha}(X)=-\sum_{x\in\mathcal{X}}P(x)\log P(x)$, collision entropy $H_{2}(X)=-\log\sum_{x\in\mathcal{X}}P(x)^{2}$, and min-entropy $H_{\infty}(X)\doteq\min_{x\in\mathcal{X}}-\log P(x)$.
Tsallis entropy [tsallis1988possible]
[…]
Onicescu informational energy [pardo1991information, calin2014informational]
is another popular measure of (un)certainty that has bred many applications in quantum mechanics, economics, ecology, social sciences, and so on. It is defined as $\rm{IE}(X):=\sum_{x\in\mathcal{X}}P(x)^{2}$ which can be connected to Rényi entropy of order-2 $H_{2}(X)=-\log\sum_{x\in\mathcal{X}}P(x)^{2}=-\log\rm{IE}(X)$, as well as the power entropy [vajda1968bounds, vajda2007generalized] $V_{2}(X)=1-\sum_{x\in\mathcal{X}}P(x)^{2}=1-\rm{IE(X)}$. There is also a concept, Onicescu’s correlation coefficient [calin2014informational], $\rho(X,Y)=\frac{\sum_{x,y}P(x)P(y)}{\sqrt{\rm{IE(X)\cdot\rm{IE(Y)}}}}$, to quantify the dependency structure between multiple variables.
These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
Appendix C Extended Context
Connection to the probabilistic Turing machines.
At some level of abstraction, one can view the LLM agents as probabilistic interactive Turing machines (PITMs) [john1977turing] that induce a distribution over action-observation transcripts. This connection may allow us to borrow some possible formal tools from the PITM literature to design UQ method for LLM agents. However, we note the following distinctions: (1) PITMs have an explicit randomness source and a defined distribution, whereas LLM agent also has a procedural uncertainty induced by the decoding time strategies; (2) while PITMs interact via fixed channels, LLM agents commonly interact with users and tools in partially observable or changing formats; and (3) the goal of LLM agents are often underspecified upfront and can be negotiated over turns depending on the situation, though PITMs have a concrete accept/output condition.
Connection to the partially observable Markov decision process.
One might draw an analogy between the agent UQ and belief tracking in Partially Observable Markov Decision Processes (POMDPs) [kaelbling1998planning]. However, the agent theme challenges the standard setups in traditional POMDPs: (1) there is no explicit belief state, but the agent carries an implicit belief in its memory; (2) actions and observations comprise language and structured strings, spanning an effectively infinite space—far from the small discrete spaces in classical POMDP; and (3) it concerns more on the agent’s uncertainty, whereas POMDP concerns mostly on environment uncertainty.
These connections highlight the distinctive edge of agent UQ, while grounding it in a well-established classical problem set.
Appendix D Note on Implementation
While we set our goal to provide a principled vision with a high-level roadmap for agent UQ, we also discuss some implementation recipes in this section to provide useful guidance for practitioners and researchers.
Action classifier.
To realize the conditional information gate, we first need to classify actions to determine whether each action will contribute to total uncertainty reduction for that task or not. As emphasized in Section 3, considering the interactivity of actions is key to accounting for agent uncertainty dynamics, as well as the evidentiality (or factuality) of the action that is solely considered for the existing approaches so far. Therefore, we may want to implement a compound classifier with both interactivity assessment and evidentiality verification. They can be implemented through a rule/syntactic-based, simple verifier [mu2024rule], an LLM judge with specialized prompting, fine-tuning [liu2025compassverifier], or a hybrid of rule-based and LLM-based [peng2025verif].
Action Category Interactivity Scenario-specific Example
Information-gathering interactive Search flight information ($\tau$-bench Airline); retrieve order status ($\tau$-bench Retail); read message box (ToolSandbox)
Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau$-bench Airline); instruct or ask the user to choose cancellation ($\tau$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
Thinking non-interactive Plan flight schedule ($\tau$-bench Airline); consider returning items ($\tau$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
State-changing tool call (write) non-interactive Update flight booking ($\tau$-bench Airline); return/cancel items ($\tau$-bench Retail); Turn on Wifi (ToolSandbox)
Providing final information to user non-interactive Summarize updated flights ($\tau$-bench Airline); summarize returned items ($\tau$-bench Retail); summarize phonebook update results (ToolSandbox)
Table 2: Example action classes in three scenarios from two benchmarks, $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
Uncertainty estimation.
In our proposed total uncertainty lower bound Eq. 6, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty. The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or the conformal prediction set size [kumar2023conformal].
However, the observation uncertainty and the initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world. If the system allows multiple trials, one can utilize a sampling-consistency-based approach to estimate the observation uncertainty, likewise the traditional black-box UQ methods. If the system does not allow this, one may try to construct a datastore that contains some relevant data from a similar task and leverage nonparameter estimation methods [ren2023outofdistribution, kotelevskii2022nonparametric]. Another line of methods can be establishing a world model to approximate the environment transition and estimate observation uncertainty from it [chae2025web, zhou2025dinowm, anonymous2026dreamphase].
Mutual information estimation.
The proposed information gating mechanism features the information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$. While the estimation of mutual information is notorious for its difficulty, recent advancements on neural estimators [belghazi2018mutual, cheng2020club, mukherjee2020ccmi, molavipour2020conditional, gritsai2025mist] shed light on this problem by enabling estimation in high-dimensional, unstructured data space.
Conditional certainty maximization process.
As an analogy to the conditional uncertainty reduction process presented in Section 4.2, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix B. Given that Onicescu’s informational energy brings its correlation coefficient as well, we may be able to implement the exact same style information-gating-based method that has the opposite direction to uncertainty reduction dynamics.
Appendix E Action Categorization
Along with the flight booking scenario in Table 1, we extend our action category analysis to two more scenarios, retail and telecommunication tasks, from $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] in Table 2.
Appendix F Missing Proof
Lemma 2 (Restatement of Lemma 1).
Let the lower bound of agent total uncertainty in Eq. 6 be $\tilde{U}(\mathcal{F}_{\leq T})$, denote $U(X):=H(X)=\mathbb{E}[-\log P(X)]$ and $\text{Info}(X;Y):=I(X;Y)=\mathbb{E}[\log\frac{P(X,Y)}{P(X)P(Y)}]$, then, we have:
$\tilde{U}(\mathcal{F}_{\leq T})$ $\geq H(E_{0},O_{0})-\sum_{i=1}^{T-1}I(O_{i},O_{0}|E_{i}\backslash O_{0})+H(A_{T}|E_{T-1},O_{T-1}),$
$\tilde{U}(\mathcal{F}_{\leq T})$ $\leq H(E_{0},O_{0})+\sum_{i=1}^{T}H(A_{i},O_{i}|E_{i-1},O_{i-1}).$
Proof.
For $i$, if $A_{i}\in\mathcal{A}^{-}$ (resp. $A_{i}\notin\mathcal{A}^{-}$), turn-level uncertainty $U(\mathcal{F}_{i})$ becomes $I(O_{i},O_{0}|E_{i}\backslash O_{0})$ (resp. $U(\mathcal{F}_{i})=U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})=H(A_{i},O_{i}|E_{i-1},O_{i-1})$). Thus, if all the intermediate actions are interactive and evidential, e.g., $A_{i}\in\mathcal{A}^{-}$ for all $1\leq i\leq T-1$ (resp. $A_{i}\notin\mathcal{A}^{-}$ for all $1\leq i\leq T-1$), our total uncertainty lower bound $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction process (resp. monotonic propagation process), deriving the above two inequalities. Here, the final termination-turn action should not be interactive, and the observation in this turn is deterministic in our considered setups (Fig. 2), resulting in the accumulation of last action uncertainty $H(A_{T}|E_{T-1},O_{T-1})$. ∎
Appendix G Discussion and Future Work
Does solely establishing the agent UQ framework actually help us to reliably infer on the agent’s performance on a task? Modern LLMs (especially after post-training) are not well-calibrated [tian2023just], so the uncertainty estimates from these ill-calibrated probabilistic models can not be directly used as a reliable performance indicator. Although we present a high-level roadmap for quantifying uncertainty for LLM agent systems, future work should also explore the calibration of LLM agents as well. The connection between calibration and accuracy [park2020calibrated, oh2024towards] of a predictor shows an exciting future work direction for joint optimization of the agent’s problem-solving capability and calibration simultaneously.
Another promising direction for future investigations would be to extend the agent UQ framework to multimodal setups such as graphical user interface agents [nguyen-etal-2025-gui], where extra challenges occur, such as many-to-many correspondence between modalities [chun2025multiplicity]. Defining and establishing a reliable UQ framework for multimodal agents would offer helpful foundations to develop robustness solutions [oh2025understanding, oh2025visual], safety [zhou2025multimodal, chen2025obvious] and more [shi2025towards].

### v2

@@
-Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents
+Uncertainty Quantification in LLM Agents: Foundations, Emerging Challenges, and Opportunities
 Abstract
@@
 We argue that UQ research must shift to realistic settings with interactive agents, and that a new principled framework for agent UQ is needed.
-This paper presents the first general formulation of agent UQ that subsumes broad classes of existing UQ setups.
-Under this formulation, we show that prior works implicitly treat LLM UQ as an uncertainty accumulation process, a viewpoint that breaks down for interactive agents in an open world.
-In contrast, we propose a novel perspective, a conditional uncertainty reduction process, that explicitly models reducible uncertainty over an agent’s trajectory by highlighting “interactivity” of actions.
-From this perspective, we outline a conceptual framework to provide actionable guidance for designing UQ in LLM agent setups.
-Finally, we conclude with practical implications of the agent UQ in frontier LLM development and domain-specific applications, as well as open remaining problems.
+This paper presents three pillars to build a solid ground for future agent UQ research: (1.
+Foundations) We present the first general formulation of agent UQ that subsumes broad classes of existing UQ setups; (2.
+Challenges) We identify four technical challenges specifically tied to agentic setups—selection of uncertainty estimator, uncertainty of heterogeneous entities, modeling uncertainty dynamics in interactive systems, and lack of fine-grained benchmarks—with numerical analysis on a real-world agent benchmark, $\tau^{2}$-bench; (3.
+Future Directions) We conclude with noting on the practical implications of agent UQ and remaining open problems as forward-looking discussion for future explorations.
 […]
@@
 Traditional LLM UQ (a) measures the uncertainty of an answer given a question, whereas the UQ for LLM reasoning (b) expands the problem by considering multi-step responses rather than a single response.
-Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau$-bench Airline; barres2025tau).
-LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent] at the next level of autonomy [Steinberger2026OpenClaw].
+Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau^{2}$-bench Airline; barres2025tau).
+LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent, anthropic2026cowork] at the next level of autonomy [Steinberger2026OpenClaw].
 In such settings, failures are no longer limited to incorrect text generation: agents may act prematurely under unresolved ambiguity, propagate errors across long interaction trajectories, or commit to outcomes that are costly or difficult to undo.
 To be deployed responsibly, agents must be able to assess and act upon the likelihood of failure [kochenderfer2015decision].
-This makes reliable uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
+This makes uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
 Despite the importance, most existing UQ research treats LLMs as static oracles: a system examined in isolation, prompted once, and evaluated by the uncertainty of a single response or a chain of responses (Figure 1, left).
 These methods implicitly assume a static system in which no new information is acquired after the initial prompt [malinin2021uncertainty, kadavath2022language, farquhar2024detecting].
-Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation.
-While appropriate for single-turn inference and non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
-Interactive agents operate in multi-turn interactive environments with feedback from users, tools, and external systems (Figure 1, right).
-For example, a flight-booking agent may need to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences.
+Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation process.
+While appropriate for single-turn inference or non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
+LLM agents in this setup operate with feedback from users and system components through multi-turn interactions to complete a complex task (Figure 1, right).
+For example, a flight-booking agent needs to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences.
 Early uncertainty should prompt information-seeking actions, while subsequent dialogue and database queries can resolve ambiguities and reduce uncertainty, enabling confident execution.
-Here, uncertainty does not strictly accumulate–it may also decrease as a direct consequence of interaction, even as the trajectory grows longer.
-In this paper, we argue that making LLM agents trustworthy requires a shift in focus: from pointwise uncertainty of answers to structured uncertainty flow in open, interactive decision processes; and a new framework is necessary for this.
-We present the following perspective:
-Uncertainty quantification for an LLM agent must incorporate interactivity to model reducible uncertainty: in non-isolated, agentic systems, total uncertainty should be allowed to decrease as new information is acquired through interaction.
-To ground this perspective in a formal discussion, we first provide a concrete definition of agent UQ and a general formulation that captures broad categories of existing UQ setups (Sec.
+This deviates significantly from the setup considered in classic LLM UQ, as the agent now spans a long interaction log updated over multiple turns that involves messages from other entities, and it also gets additional information from environments.
+This gap raises questions about the applicability of existing UQ methods to agentic setups.
+In this paper, we argue that building a trustworthy LLM-driven agentic system requires a shift in focus: from pointwise uncertainty of the final answer to structured uncertainty dynamics in an open, interactive decision process, and a new framework is necessary for this.
+To initiate a formal discussion, we first provide a concrete definition and a general formulation of agent UQ that captures broad categories of existing UQ setups (Sec.
 3).
 Specifically, we model the agent’s problem-solving trajectory as a stochastic process over actions, observations, and environment states, represented as a simple graphical model.
-Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ approaches arise as special cases.
-This unified view also exposes a fundamental limitation of existing methods: uncertainty is implicitly treated as accumulating monotonically along the trajectory, regardless of the role played by different actions; thus fails to relate uncertainty to a task reward.
-To address this limitation, we introduce in Sec.
-4 a conceptual framework that models agent UQ as a conditional uncertainty reduction process, through an information-gating mechanism that operationalizes the distinction between uncertainty-increasing and uncertainty-reducing actions by considering their interactivity.
-We further derive the analytic bounds that improve the interpretability of the resulting uncertainty estimates.
-In addition, Sec.
+Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ problems can be cast as special cases.
+Based on this foundation, we identify four technical challenges in agent UQ that call for actions from communities.
+(1) Selection of uncertainty estimator: while it is well-known that existing uncertainty estimators have their own weaknesses, agentic scaffolding amplifies these into severe limitations or hard constraints; (2) Uncertainty of heterogeneous entities: how can we derive estimates over the content generated by external entities having different underlying distributions?; (3) Modeling uncertainty dynamics in interactive systems: traditional weighted-average-style uncertainty aggregation may not be suitable to modeling uncertainty dynamics in the non-isolated, open environment; (4) Lack of fine-grained benchmarks: our pilot survey reveals that granular benchmarks, such as turn-level evaluation, are scarce for LLM agents.
+Then, Sec.
 5 discusses practical implications of agent UQ across diverse domains, including healthcare, software engineering, and robotics, as well as frontier LLM research on adaptive reasoning and post-training.
-Finally, we outline open problems unique to uncertainty modeling in agentic systems (Sec.
+Finally, we outline open problems specific to uncertainty modeling in agentic systems (Sec.
 6).
-Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware agentic systems.
-  • We formalize agent UQ as a distinct problem setup and show why existing LLM UQ formulations are insufficient for interactive agent setups.
-  • To tackle that limitation, we present a conceptual framework for agent UQ by viewing it as a conditional uncertainty reduction process.
-  • To give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
+Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware, reliable agentic systems.
+  • We cast agent UQ as a distinct problem setup, and build a foundation for that by providing a concrete definition and a general formulation that expresses many existing UQ setups as special cases.
+  • We crystallize the emerging challenges in agent UQ with numerical studies on a real-world benchmark $\tau^{2}$-bench [barres2025tau] with GPT-4.1 [openai2025gpt41] and Kimi-K2.5 [kimi2026].
+  • To further give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
 2 Related Work
 UQ in LLMs.
-In contrast to classic UQ [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs.
-Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over samples [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal].
+In contrast to classic UQ in machine learning [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs.
+Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over multiple generations [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal].
 While promising, they mostly use single-step question-answering as a testbed, which limits their practical applicability to agentic setups.
@@
 This simple abstraction describes some representative agentic prompting methods such as ReAct [yao2022react].
-See Appendix A.1 for details.
+See Appendix B.1 for details.
 UQ in reasoning and agent setups.
 Beyond single-step generation, researchers show growing interest in measuring the confidence of LLMs during their reasoning.
-This kind of (un)certainty information has been used to improve the quality of response [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning].
-Recently, there have also been attempts to model uncertainty in interactive and/or agentic setups [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal].
-Although UQ here aims to model uncertainty dynamics over long-horizon generation, existing methods do not capture the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new framework for agent UQ.
+This (un)certainty information has been used to improve response quality [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning].
+Recently, there have been attempts to model uncertainty in interactive or agentic inference [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal, zhang2026confidence, zhang2026auq].
+Although aiming to model uncertainty dynamics over long-horizon generation, they do not reflect uncertainty from different entities (e.g., users) nor concern the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new UQ framework under agentic scaffolding.
 3 A General Formulation for Agent UQ
@@
 The observation $O$, on the other hand, can be a database status report, tool-calling results, or the user’s clarification.
-Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., code.
+Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., programming language.
 Concretely, given a task specification $E_{0}$, e.g., a flight booking, a user makes a stochastic initial query $O_{0}$, and the agent first requests the user’s ID to get relevant information.
@@
 Definition 1 (Stochastic Agent System).
-Let $E_{i}$ be an environment state, a mixture of a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and a state of the system database, which is not directly accessible to the agent (only partially observable via tool-calling).
+Let $E_{i}$ be an environment state, a mixture of (1) a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and (2) a state of the system database, which is not directly accessible to the agent—only partially observable via tool-calling.
 Let $O_{i}$ and $A_{i}$ be the $i$-th turn observation and action derived from distributions $P$ and $P_{\pi,\mathcal{T}}$, respectively, where $\mathcal{T}$ indicates a tool set and $\pi$ is an LLM policy.
 Then, under a task specification $E_{0}$ and the user’s initial query $O_{0}$, we define the trajectory as a sequence of turns $\mathcal{F}_{\leq T}=\{(A_{t},E_{t},O_{t})\}_{t=0}^{T}$ where $\mathcal{F}_{0}=(E_{0},O_{0})$.
-Given a deterministic update function $h(\cdot)$, the generative process for each step $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
+Given a deterministic update function $h(\cdot)$, the generative process for each turn $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
 To express this trajectory with a simple mathematical representation, we introduce a dynamic Bayesian network [murphy2002dynamic] in Figure 2.
 With this graphical model, dependency structures of $A$, $E$, and $O$ can be explained as follows: (1) since the environment state $E_{i-1}$ contains the entire conversation history of $A_{\leq i-1}$ and $O_{<i-1}$ up to the i1i-1 step [^1] An adaptive memory structure [xu2025mem, xiong2025memory] can replace this ever-growing memory., the current action $A_{i}$ depends only on $E_{i-1}$ and $O_{i-1}$; (2) the current observation $O_{i}$ from the environment also depends only on $A_{i}$ and $E_{i}$; (3) the transition of the environment state ($E_{i-1}\rightarrow E_{i}$) can be defined as a deterministic function [^2] In general, this system transition can be also stochastic, but we consider a simplified yet still realistic scenario.
-$h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system state given the current action.
+$h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system database state given the current action.
 Under this Bayesian network, the joint probability for the trajectory can be factorized as below,
@@
 In the machine learning field, the most popular choices of the uncertainty function $U(X)$ are Shannon entropy [shannon1948mathematical] $\mathbb{E}[-\log P(X)]$, negative log probability $\log P(X=x)$, and their variants.
-These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suit for sequential modeling.
-For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix A and B for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
+These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suitable for sequential modeling.
+For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix B and C for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
 $U(\mathcal{F}_{\leq T})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$
-$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$
-where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ denotes an initial query uncertainty plus a task volatility.
+$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$ (1)
+where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ is an initial query uncertainty plus a task specification volatility.
 Desideratum.
@@
 where $r(\cdot)$ denotes a real-value reward function commonly set as a success-failure binary annotator.
-We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus.
-See Appendix C for discussion.
+We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded by, while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus.
+See Appendix D for the relevant discussion.
 3.2 A Unified View on Existing UQ Setups
@@
 3 to Eq.
-3, while a few consider it [leang2025picsar].
+4, while a few consider it [leang2025picsar].
 Meanwhile, rather than summing up all the step-wise uncertainties, some approaches consider the most uncertain one, i.e., the lowest confidence used in fu2025deep, which can be represented by Eq.
-4.
+5.
 Finally, Eq.
-5 expresses various weighted average methods, such as a standard length-normalized uncertainty [kang2025scalable] by setting $w_{i}=\frac{1}{T}$, and a tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies, i.e., top-K mean [fu2025deep] or model-based estimates [zhao2025uncertainty].
+6 expresses various weighted average methods, e.g., length-normalized uncertainty [kang2025scalable] with $w_{i}=\frac{1}{T}$, and tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies [zhao2025uncertainty].
 Connection to process reward modeling.
-We further draw a connection between this trajectory-level UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward.
+We further draw a connection between this multi-turn UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward.
 For example, lightman2024lets proposed the product of the step-wise probabilities and the minimum probability between steps as the aggregation method.
@@
 4 and Eq.
-4, respectively.
-Figure 3: Limitation of existing UQ (a) and our suggestion (b).
-Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure.
-We urge expanding a dimension of interest, interactivity, and moving on to reducible uncertainty modeling of agents.
-Limitation of existing frameworks.
-From this unified view, we see that they all model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn.
-One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa.
-However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward.
-See Fig.
-3 (a), a scenario wherein a user does not fully specify their goal upfront.
-After searching for options, the agent either reasons itself or interacts with the user to delegate decisions.
-Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence.
-Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 1).
-4 How to Model Uncertainty Dynamics of Interactive Agents?
-4.1 Unique Challenge
-In contrast to the LLM UQ, agent UQ spans a multi-turn trajectory in an open environment.
-This presents a unique challenge: modeling cascade uncertainty in an interactive system.
-Here, the desired UQ method should be able to precisely aggregate uncertainties across the entire trajectory and account for additional information gained through interactions.
-While there is emerging research on step-wise uncertainty aggregation [kang2025scalable, fu2025deep], they examine an isolated, noninteractive setting where an agent does not receive any external feedback during inference (e.g., Fig.
-1, b).
-Meanwhile, agent UQ considers a setup full of interactions with a user and an environment, making it a non-isolated, interactive system.
-Since an agent in this open world can gather additional information, make clarifications, and request confirmations at each step by interacting with the user and external databases through conversations and tool calls, the uncertainty dynamics in this interactive system includes not only propagations of turn-level uncertainty, but also reductions.
-As shown in Fig.
-3, the existing perspective does not consider “interactivity” in agent UQ and prohibits reducible uncertainty modeling.
-In this paper, we argue that we must introduce interactivity in agent UQ to model bi-directional uncertainty flow.
-4.2 Agent UQ as a Conditional Uncertainty Reduction Process
-We need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, depending on the agent’s interactivity at each step.
-While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec.
-3.
-Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
-$U(\mathcal{F}_{\leq{T}})=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})]$
-$=U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})[1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}]$
-$\geq U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})g(\mathcal{F}_{i}).$ (6)
-Here, we introduce a conditional gating function $g(\cdot)$ that reduces or increases a current turn uncertainty depending on the action.
-That is, for a set of valid uncertainty reduction actions, $\mathcal{A}^{-}$, we define:
-$g(\mathcal{F}_{i})=\begin{cases}\frac{-\text{Info}(O_{i};O_{0}|E_{i}\backslash O_{0})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{if}\penalty 10000\ A_{i}\in\mathcal{A}^{-},\\ 1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{otherwise.}\end{cases}$
-where $\text{Info}(\cdot;\cdot)$ denotes mutual information or pointwise mutual information that measures the amount of information gain by having the current observation $O_{i}$ w.r.t.
-initial query $O_{0}$ given conversation history $E_{i}\backslash O_{0}$.
-For each turn, if the agent’s action is classified as an interactive (inviting a user or a tool) and evidential one (factual or not in conflict with stored data), then $g(\cdot)$ produces the amount of information gain with negative sign to realize uncertainty reduction; otherwise, it propagates uncertainty (See Fig.
-4).
-Figure 4: Illustration on the proposed agent UQ paradigm. We propose the conditional uncertainty reduction process for LLM agents by discerning interactive and evidential actions from others.
-Importantly, the signed gating function handles uncertainty in a more interpretable manner than traditional estimates by letting not only quantitative but also directional interpretations.
-Besides, this uncertainty model allows us to derive the analytic extrema, i.e., maximum and minimum, of total uncertainty in closed forms as shown in Lemma 1.
-Action Category Interactivity Example
-Information-gathering interactive Agent uses a read tool (e.g., get_reservation_details or search_flights) to retrieve data or asks the user for missing information, e.g., “Could you provide your reservation number and last name?”.
-Asking clarification or confirmation to user interactive Agent asks the user preferences or confirmation on decision, such as, “Do you want me to proceed with booking FQ8APE?”.
-Thinking non-interactive Agent plans the future action sequence based on the previous trajectory.
-State-changing tool call (writing) non-interactive Agent calls the tools that modify the database (e.g., cancel_reservation, book_reservation, update_reservation_flights) commit to an outcome.
-Providing final information to user non-interactive Agent reports the result of an action, such as “Your reservation has been cancelled; your refund will be processed” or “Your flight has been rebooked to SFO departing at 8 a.m.”.
-Table 1: Example action classes in Airline booking scenario in $\tau$-bench.
-A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty.
-See Appendix E for extended analyses.
-Lemma 1 (Extrema of Information Gating).
-Let the lower bound of agent total uncertainty in Eq.
-6 be $\tilde{U}(\mathcal{F}_{\leq T})$, denote $U(X):=H(X)=\mathbb{E}[-\log P(X)]$ and $\text{Info}(X;Y):=I(X;Y)=\mathbb{E}[\log\frac{P(X,Y)}{P(X)P(Y)}]$, then, we have:
-$\tilde{U}(\mathcal{F}_{\leq T})$ $\geq H(E_{0},O_{0})-\sum_{i=1}^{T-1}I(O_{i},O_{0}|E_{i}\backslash O_{0})+H(A_{T}|E_{T-1},O_{T-1}),$
-$\tilde{U}(\mathcal{F}_{\leq T})$ $\leq H(E_{0},O_{0})+\sum_{i=1}^{T}H(A_{i},O_{i}|E_{i-1},O_{i-1}).$
-Proof.
-If $A_{i}\in\mathcal{A}^{-}$ (resp.
-$A_{i}\notin\mathcal{A}^{-}$) for all $1\leq i\leq T-1$, then $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction (resp.
-propagation) process, deriving above inequalities (See Appendix F for details).
-∎
-Interpretation.
-Lemma 1 states that an agent’s total uncertainty can be expressed as an ever-decreasing process starting from a task-specific initial uncertainty reduced by the amount of extra information delivered by current observation, given the previous trajectory till its final behavior in the ideal case; whereas it is continually stacked up by the amount of joint entropy between action and observation per turn in worst case.
-Together with the directional interpretation turn-level uncertainty (increase or decrease), these analytic extrema help us better understand the numerical uncertainty, which is usually hard to interpret [van2019communicating, belem2024perceptions, kirchhof2025self].
-Sketch of implementation.
-To operationalize this idea, we first need an action classifier to determine whether $A_{i}\in\mathcal{A}^{-}$ or not.
-One can adopt a model-based evaluator [kim2025beyond] and a rule-based verifier [zeng2025reinforcing] to assess the interactivity and evidentiality of actions, respectively, where $A_{i}\in\mathcal{A}^{-}$ if and only if $A_{i}$ is both evidential and interactive.
-An example categorization of actions in an airline assistant task is provided in Table 1.
-After that, four different quantities need to be estimated: (1) initial query uncertainty, (2) action uncertainty, (3) observation uncertainty, and (4) mutual information between the query and observation given the previous trajectory.
-For these estimations, we can leverage the output probability [malinin2021uncertainty, aichberger2024rethinking], consistency in multiple trials [farquhar2024detecting], world modeling [chae2025web, anonymous2026dreamphase], and a learning neural estimator [molavipour2020conditional, gritsai2025mist].
-We leave the actual implementation for future work.
-See Appendix D for additional notes.
+5, respectively.
+4 Technical Challenges in Agent UQ
+We discuss why agentic setups are more challenging for UQ than non-agentic ones, identifying four distinct challenges.
+To motivate discussion, we present preliminary results on $\tau^{2}$-bench using existing UQ methods: negative log-likelihood (NLL), entropy over token probabilities (Entropy), and verbalized confidence, each aggregated over all tokens or turns generated by an agent throughout the trajectory (Appendix F has details).
+4.1 Selection of Uncertainty Estimator
+Approach Access Inference Overhead Theoretical Ground
+Probability-based Free
+Consistency-based Significant
+Verbalized Confidence Neglectable
+Table 1: Comparison between UQ approaches. Existing UQ approaches have unique weaknesses that become severe obstacles in agentic setups.
+So what kind of uncertainty estimator should we adopt to model the agent’s uncertainty, $U(\cdot)$ in Eq.
+3.1?
+We can start by examining three representative approaches that have been advanced in LLM UQ literature so far: output probability-based (and/or internal hidden state-based), response consistency-based, and prompting-based verbalized confidence methods.
+These approaches involve clear trade-offs along three dimensions: accessibility, inference-time cost overhead, and theoretical grounding.
+Probability-based methods exhibit limited applicability due to their reliance on access to output probabilities.
+While consistency-based and verbalized confidence methods do not face this accessibility constraint, they incur substantial inference-time overhead from repeated generation and lack theoretical grounding, respectively.
+Scenario Avg. $r$ NLL Entropy Verb. Conf.
+AUROC $\rho$ $\tau$ AUROC $\rho$ $\tau$ AUROC $\rho$ $\tau$
+GPT-4.1 on Retail 0.509 0.597 0.169 0.138 0.580 0.139 0.114 0.575 0.179 0.170
+GPT-4.1 on Telecom 0.517 0.624 0.214 0.176 0.611 0.192 0.158 0.685 0.330∗∗∗ 0.286∗∗∗
+Kimi-K2.5 on Retail 0.447 0.469 -0.054 -0.044 0.468 -0.056 -0.045 0.523 0.039 0.032
+Kimi-K2.5 on Telecom 0.965 0.645 0.093 0.076 0.664 0.104 0.086 0.580 0.051 0.042
+Table 2: Evaluation of uncertainty estimators on $\tau^{2}$-bench.
+NLL and Entropy predict task failure (1r1-r); Verbalized confidence predicts task success ($r$), where $r$ denotes the reward; $\rho$ and $\tau$ denote Spearman and Kendall rank correlation coefficient, respectively.
+The significance is noted as $^{*}p<0.05$ and $^{***}p<0.001$.
+Although this analysis applies to the classical single-turn QA setup as well, agentic scaffold turns these weaknesses into much more serious challenges: (1) probability-based methods cannot be applied to most frontier LLMs that centered at challenging agent benchmarks, as they do not consistently provide probability outputs; moreover, free-form long generation makes aggregated token probabilities less informative; (2) consistency-based methods become infeasible due to their prohibitively high inference cost in long-horizon, multi-turn settings; and (3) dynamically expanding context memory of agent (i.e., $E_{i}$ in Eq.
+3.1), often coupled with noisy observations, results in increasingly inflated and unreliable verbalized confidence.
+A pilot study in Table 2 provides some results of probability-based and verbalized confidence methods for estimating the trajectory uncertainty $U(\mathcal{F}_{\leq T})$ with aggregated action uncertainty $\frac{1}{T}\sum_{i=1}^{T}U(A_{i})$.
+In many cases, the considered methods show close-to-random classifier performance in predicting the agent’s success/failure.
+We believe that establishing a theoretically grounded verbalized confidence method is a promising future direction, given its vivid merits in agentic setups such as accessibility and minimal extra inference cost.
+4.2 Uncertainty of Heterogeneous Entities
+Figure 3: Observation uncertainty on $\tau^{2}$-bench.
+We compare distributions over observation uncertainty estimates, i.e., average NLL over all user messages, from the agent LLM and the ground truth user simulator LLM.
+Inference of agents involves not only the actions $A_{i}$ from the agent itself, but more importantly, observations $O_{i}$ through the conversations with the user and interactions with the environment via tools.
+There are no trivial means to estimate the uncertainty of observations that come from these heterogeneous external entities, departing from the epistemic uncertainty of the agent-self.
+One may argue that we can just feed the whole trajectory, including the observation strings from the user and tools, into the agent and then use its output probability $P_{\pi,\mathcal{T}}(O_{i}|A_{i},E_{i})$ to approximate the underlying true observation distribution $P(O_{i}|A_{i},E_{i})$ per turn and corresponding uncertainty $U(O_{i}|A_{i},E_{i})$.
+However, the distribution of human language is definitely different from that of LLMs [munozortiz2024contrasting, alex2025do], and the distribution over the tool-calling outcome strings might be different too.
+To quantify this gap, we conduct a preliminary experiment on $\tau^{2}$-bench, where the user is simulated with another LLM equipped with a distinct prompt from that of the agent LLM.
+In Figure 3, we compare the trajectory-level average NLL density plots driven from the ground truth user simulator and the agent LLM.
+We see remarkable deviations between them, stressing the need for deliberation to estimate observation uncertainty.
+Adopting auxiliary LLM as an approximated world model can be a possible solution, where we see some promises in Figure 8.
+4.3 Modeling Uncertainty Dynamics in Interactive Systems
+One of the most distinctive features of agent UQ is that the model now operates in a non-isolated environment to generate a trajectory from iterative interactions.
+This presents a unique challenge: modeling multi-turn cascade uncertainty in an interactive system.
+Here, the desired UQ method should not only model the propagation of uncertainty across the trajectory but also account for additional information gained through interactions.
+To be specific, an agent in this open environment can gather extra information, make clarifications, and request confirmations at each step by interacting with the user and databases through conversations and tool-calling.
+These interactive actions are the driving force of goal achievement in the complex (and usually under-specified) agentic tasks in the wild [li2026thinking].
+Therefore, the uncertainty model, which can reliably inform the agent’s failures in this setting, should reflect conditional uncertainty reduction through information-seeking behavior as well as the uncertainty propagation.
+Figure 4: Uncertainty evolution across the trajectory on $\tau^{2}$-bench.
+We visualize the evolution of action uncertainty estimates across the normalized turn index, averaged over the entire trajectories from two groups: task success and failure.
+Unfortunately, most existing multi-step UQ approaches discussed from Eq.
+3 to Eq.
+6 do not consider the interactivity of actions for each turn, and only focus on uncertainty propagation, while neglecting the reducible nature of uncertainty dynamics in non-isolated, interactive systems.
+In Figure 4, we present the uncertainty evolution of a naive averaging-based uncertainty aggregation method (Eq.
+6), which is averaged over all trajectories from both success and failure groups.
+This simple, action-independent aggregation fails to meaningfully discern the failure group from the success group; Even the failure group achieves a sharper uncertainty decrease in the later part of trajectories in Telecom domain.
+This implies that the action-independent naive cascade modeling cannot robustly identify the agent’s failure both in the process and at the termination of the task, indicating the necessity of a new framework, pursuing explicit action-conditional modeling.
+We discuss a possible recommendation on this action-conditional uncertainty dynamics model in Appendix LABEL:sec:adpx:action-conditional.
+4.4 Lack of Fine-grained Benchmarks
+In the meantime, similar to the process reward modeling [choudhury2025process], annotating long-horizon agent trajectories incurs non-trivial effort and cost.
+The absence of fine-grained benchmarks becomes a natural bottleneck to developing a solid UQ method for agents.
+In this subsection, we provide a mini-survey of existing agent or multi-turn interaction benchmarks (See Appendix G), focusing primarily on LLM-centric agents rather than multimodal or physical ones.
+Specifically, we categorize 44 benchmarks into three classes depending on evaluation granularity: trajectory-level (conducted once at the end of the trajectory), milestone-level (involving several intermediate milestones or events), and turn-level (conducted at every single turn).
+Figure 5 reveals that only 4 out of 44 benchmarks provide turn-level annotation, which is very scarce compared to the trajectory-level benchmarks.
+Besides, although the milestone-level alternative shows better population than the turn-level one, it is still lacking, resulting in three times less than the trajectory-level one.
+Figure 5: Survey on 44 modern agent benchmarks. The number denotes the percentage of papers in each category.
+There are several obstacles to developing fine-grained benchmarks for agentic inference setups.
+First, agent benchmarks often involve very long rollouts that can easily exceed a hundred steps, making it hard to manually annotate each step.
+Although using frontier models as a substitute for human labeling has become increasingly common, running such models on long-horizon tasks can be prohibitively expensive for many researchers.
+Moreover, agent benchmarks are usually defined by complex policies; even for humans, it may take a couple of hours to fully understand the task configuration.
+To accelerate the development of agent UQ, we should pay greater attention to scalable dataset construction strategies for agentic inference.
 5 Practical Implications
-Developing an agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments.
-We outline implications for frontier LLM research and three specialized domains: healthcare, coding, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
+Developing the agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments.
+We outline implications for frontier LLM research and three specialized domains: healthcare, programming, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
 5.1 Advancing Frontier LLMs
@@
 5.3 Reliable Software Engineering Agents
-Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for an agent UQ framework.
-Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, or commit to a fix.
+Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for a reliable agent UQ framework.
+Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, commit to a fix, or interact with the user.
 Besides, considering uncertainty-triggered rollback and branching algorithm, in which an agent treats increasing uncertainty after an action as a signal to revert or branch by mirroring the checkpoint-rollback workflows of human engineers.
 These highlight the potential of UQ for reliable agentic coders.
-5.4 Relible Embodied Agents in Cyber-Physical Systems
+5.4 Reliable Embodied Agents in Cyber-Physical Systems
 Embodied agents in cyber-physical systems face heterogeneous uncertainty from sensing, dynamics, and human intent [li2024embodied, fung2025embodied, team2025gemini], and they actively reduce uncertainty through interaction by re-sensing the environment or interacting with humans before committing to irreversible physical actions [xu2025compliant].
@@
 6 Open Problems
-Although we have so far presented a conceptual roadmap and the practical implications of agent UQ, significant open problems remain as follows.
-Developing rigid benchmarks at scale.
-As in the case of process reward modeling mentioned in Section 3.2, annotating across long-horizon trajectories introduces non-trivial efforts and costs [choudhury2025process].
-Moreover, as true uncertainty is unmeasurable, i.e., only a proxy-based evaluation is possible, establishing benchmarks for agent UQ stands on the next level of difficulty than the static LLM evaluation [ielanskyi2025addressing].
-Investigation on defining proper agent performance measures [kapoor2025holistic, hendrycks2025definition] with scalable evaluation and annotation methods for agent rollouts [yang2024react, chang2024agentboard, kim2025beyond] would be a crucial stepping stone.
-Long-horizon black-box uncertainty estimation.
-LLM agents are mostly offered by proprietary services, where we cannot access the output probability of the agent.
-While model owners and providers can leverage low-cost white-box UQ approaches, others may have to rely on expensive black-box UQ that requires heavy sampling per action over a long trajectory [zhang2024luq].
-This poses a great challenge for low-resource users in developing UQ with state-of-the-art agents.
-Solution multiplicity.
+We have so far presented a roadmap for agent UQ with a general formulation, identified challenges, and practical implications, but important open problems remain as follows.
+Intrinsic solution multiplicity.
 Is high uncertainty per step due to the agent’s lack of knowledge about the right action for that step?
@@
 Thus, each intermediate step is intrinsically ambiguous and allows multiple valid actions, making it impossible to specify the source of the high uncertainty.
-UQ in multi-agent or self-evolving agent systems.
-While we confine our setup to single-agent, non-evolving agent systems, there has been remarkable advancement in multi-agent systems [guo2024large, wu2024autogen] and self-evolving agents [wang2024voyager, jiang2025adaptation].
-Modeling the joint uncertainty dynamics of multiple agents or considering the non-stationary toolkit and agent parameters over time brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
+Graph-based parallelism/multiplicity modeling might be worth exploring as a possible solution here [wu2025regular, choi2026modex].
+Evaluation beyond task failure.
+Evaluation of predictive uncertainty has been conducted by measuring the informativeness of uncertainty estimates when predicting the incorrectness of the final answer.
+While suitable in single-turn QA settings, it becomes insufficient when we move to agentic LLM setups.
+In such setups, the sole prediction of eventual task failure collapses a rich, structured problem-solving process into a single scalar judgment, discarding useful information that makes uncertainty more actionable.
+We envision a new evaluation protocol that integrates dynamic solution multiplicity, task difficulty, and irreducible external ambiguity, as well as task failure, that may redefine uncertainty from a multi-facet view [guo2025uncertaintyprofiles].
+This reframing calls for new metrics that credit not just predictive correlation with outcomes but the informativeness of uncertainty signals in multiple factors in agentic scaffolding, which is yet to be explored.
+Uncertainty modeling in multi-agent systems.
+Although we confine our attention to a single-agent environment, there has been notable advancement in multi-agent systems (MAS) [Liang2023EncouragingDT, du2024improving, guo2024large, wu2024autogen, choi2025debatevote].
+Uncertainty dynamics in this MAS can be leveraged to improve the quality of inter-agent communication [yoffe2025debunc] and mitigate debate collapse [tang2026value], thus worth exploring.
+Unfortunately, while there are some works that borrow the strength of collective intelligence from MAS to improve UQ quality in a single-turn QA setup [yang2024confidence, feng2025rethinking], no existing work in this field attempts to design a UQ method tailored for MAS.
+Modeling the joint uncertainty dynamics of multiple agents brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
+Uncertainty modeling in self-improving agents.
+In our setup, we assume that the agent only operates within a single episode.
+However, the agent research community has recently shown an increasing interest in self-evolving/improving agents [bai2022constitutional, shinn2023reflexion, NEURIPS2023_91edff07, wang2024voyager, jiang2025adaptation] that adapt across multiple episodes.
+Uncertainty information can also be utilized in this setup for reliable continual adaptation [zhang2026selaur].
+However, the interactive self-evolving environment adds another non-trivial factor that shifts uncertainty dynamics as shown in huang2025beyond.
+Future exploration in this setup should reflect the non-stationary toolkit, context memory, and agent model parameters across multiple episodes.
 7 Conclusion
-We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the bi-directional dynamics model of uncertainty.
+We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the sequential dynamics model of uncertainty.
 We laid the foundation for this by providing a concrete definition and general formulation of agent UQ that abstracts a wide variety of existing UQ approaches.
-To tackle the limitations of existing framing, we presented a novel perspective, conditional uncertainty reduction process, with actionable guidelines and a theoretical analysis.
-As a concluding remark, we quote Schrödinger’s note on living in physics [schrodinger1974what]: “Every living means an increase in entropy of the part of the world, and it can only survive by continually drawing negative entropy from its environment.”
+With that, we identified four key challenges emerging in UQ on agentic scaffolding with detailed explanations and numerical analyses.
+We further note the future direction to explore agent UQ from the perspectives of real-world applications and open problems to provoke forward-looking discussions.
+We hope the three pillars provided in this work help communities design their upcoming works, products, and more.
 Limitations
-Although this paper provided a novel high-level vision for agent UQ with a crystallized formulation, we built this vision on a relatively simple setup with some assumptions: reliable observation $O_{i}$ from the environment and user, and deterministic environment update $E_{i}\rightarrow E_{i+1}$.
-However, agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment, where our current proposal does not ensure its effectiveness.
+Agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment.
+However, we do not dive deeper on that settings, and our formulation might not provide a proper abstraction.
 In that scenario, we may need to explicitly model uncertainty in the agent’s evidence stream as well, e.g., trust/ignorance about tool outputs, retrieval, and user-provided facts, by leveraging formalisms, such as subjective logic [jsang2018subjective, cheng2020there], imprecise probability theory [walley1991statistical], and formal belief representation [formalbelief22], to safely manage memory and act under unreliable observations in stochastic world.
@@
 Acknowledgments
-We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, and Kyungwoo Song for their sharp feedback on the draft.
+We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, Hoyoon Byun, and Kyungwoo Song for their sharp feedback on the draft.
 This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23- 1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P.
@@
 Appendix
-We cover the following content in this Appendix.
-  • (A) Additional Details on Formulation
-  • (B) Alternative Uncertainty Measures
-  • (C) Extended Context
-  • (D) Note on Implementation
-  • (E) Action Categorization
-  • (F) Missing Proof
-  • (G) Discussion and Future Work
-Appendix A Additional Details on Formulation
-A.1 Expression of various agentic prompting methods under our formulation.
+Appendix A Revision History
+A prototype version of this paper titled "Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents" was released on arXiv in early February 2026.
+It was specifically targeted at the third challenge described in the current Section 4.3 by proposing a conceptual framework, conditional uncertainty reduction process, to enable action-conditional uncertainty dynamics modeling for interactive agents, which is available now at Section E in this Appendix.
+After receiving some early feedback from the community, we revised this article to cover a broad range of practical challenges emerging under agentic scaffolding, supported by quantitative analyses, rather than focusing narrowly on a single challenge and its corresponding solution.
+In the main body of the paper, Section 4 has undergone a major revision, while the remaining sections were subject to only moderate or minor edits.
+Appendix B Additional Details on Formulation
+B.1 Expression of Various Agentic Prompting under Our Formulation
 In Section 3.1, we have introduced a definition of stochastic agent system (Def.
-1) with a probabilistic graphical model of specific conditional dependency assumptions 2.
+1) with a probabilistic graphical model of specific conditional dependency assumptions in Figure 2.
 Under this model, we can abstract some representative agentic prompting methods [ahn2022can, yao2022react, shinn2023reflexion, patil2025berkeley] with a unified language yet distinctive conditions as below.
@@
 Overall, this suggests that our agent formulation is general and flexible enough to capture a broad class of representative agentic prompting scenarios.
-A.2 Uncertainty Instantiations and Total Uncertainty Expansion.
+B.2 Uncertainty Instantiations and Total Uncertainty Expansion
 After defining the agent UQ in Def 2, we presented a trajectory-level total uncertainty that expresses the joint uncertainty across multiple turns in an additive form of the initial query uncertainty, action uncertainty, and observation uncertainty.
-Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
+Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i})$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
   1. 1.
@@
 which directly drives our uncertainty expansion for $U(\mathcal{F}_{\leq T})$ in Section 3.1.
-Appendix B Alternative Uncertainty Measures
+Appendix C Alternative Uncertainty Measures
 Although we have mainly focused on the above three information-theoretic uncertainty measures throughout the paper, there are plenty of alternatives one can consider depending on the problem setups.
 In this section, we examine some possible alternative uncertainty measures, Rényi entropy, Tsallis entropy, and informational energy.
+These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
 Rényi entropy [renyi1961measures]
@@
 There is also a concept, Onicescu’s correlation coefficient [calin2014informational], $\rho(X,Y)=\frac{\sum_{x,y}P(x)P(y)}{\sqrt{\rm{IE(X)\cdot\rm{IE(Y)}}}}$, to quantify the dependency structure between multiple variables.
-These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
-Appendix C Extended Context
+Appendix D Extended Context
 Connection to the probabilistic Turing machines.
@@
 These connections highlight the distinctive edge of agent UQ, while grounding it in a well-established classical problem set.
-Appendix D Note on Implementation
-While we set our goal to provide a principled vision with a high-level roadmap for agent UQ, we also discuss some implementation recipes in this section to provide useful guidance for practitioners and researchers.
+Appendix E Action-Conditional Uncertainty Dynamics Model
+Figure 6: Limitation of existing UQ (a) and our suggestion (b).
+Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure.
+We recommend expanding a dimension of interest, interactivity, and moving on to tackle the third challenge (Sec, 4.3) by allowing reducible uncertainty modeling of agents.
+Motivation.
+From a unified view discussed in Section 3.2, we see that most of the existing UQ methods model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn.
+One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa.
+However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward.
+See Fig.
+6 (a), a scenario wherein a user does not fully specify their goal upfront.
+After searching for options, the agent either reasons itself or interacts with the user to delegate decisions.
+Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence.
+Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 2).
+From our pilot experiments (Figure 9 and Figure 10 provided in Appendix F.2), we observed that an action-independent naive aggregation strategy could not faithfully separate between the success and failure groups on average, resulting in full overall performance reported in Table 2, regardless of the types of uncertainty estimator.
+Recommendation: conditional uncertainty reduction process.
+We therefore need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, explicitly depending on the agent’s interactivity at each step.
+While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec.
+3.
+Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
+$U(\mathcal{F}_{\leq{T}})=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})]$
+$=U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})[1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}]$
+$\geq U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})g(\mathcal{F}_{i}).$ (7)
+Here, we introduce a conditional gating function $g(\cdot)$ that reduces or increases a current turn uncertainty depending on the action.
+That is, for a set of valid uncertainty reduction actions, $\mathcal{A}^{-}$, we define:
+$g(\mathcal{F}_{i})=\begin{cases}\frac{-\text{Info}(O_{i};O_{0}|E_{i}\backslash O_{0})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{if}\penalty 10000\ A_{i}\in\mathcal{A}^{-},\\ 1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{otherwise.}\end{cases}$
+where $\text{Info}(\cdot;\cdot)$ denotes mutual information or pointwise mutual information that measures the amount of information gain by having the current observation $O_{i}$ w.r.t.
+initial query $O_{0}$ given conversation history $E_{i}\backslash O_{0}$.
+For each turn, if the agent’s action is classified as an interactive (inviting a user or a tool) and evidential one (factual or not in conflict with stored data), then $g(\cdot)$ produces the amount of information gain with negative sign to realize uncertainty reduction; otherwise, it propagates uncertainty (See Fig.
+7).
+Figure 7: Illustration on the proposed agent UQ paradigm. We propose the conditional uncertainty reduction process for LLM agents by discerning interactive and evidential actions from others.
+Importantly, the signed gating function handles uncertainty in a more interpretable manner than traditional estimates by letting not only quantitative but also directional interpretations.
+Besides, this uncertainty model allows us to derive the analytic extrema, i.e., maximum and minimum, of total uncertainty in closed forms as shown in Lemma 1.
+E.1 Implementation Sketch: Conditional Uncertainty Reduction Process
+We discuss a possible implementation recipe for this conditional uncertainty reduction process to provide useful guidance for practitioners and researchers.
 Action classifier.
@@
 They can be implemented through a rule/syntactic-based, simple verifier [mu2024rule], an LLM judge with specialized prompting, fine-tuning [liu2025compassverifier], or a hybrid of rule-based and LLM-based [peng2025verif].
-Action Category Interactivity Scenario-specific Example
-Information-gathering interactive Search flight information ($\tau$-bench Airline); retrieve order status ($\tau$-bench Retail); read message box (ToolSandbox)
-Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau$-bench Airline); instruct or ask the user to choose cancellation ($\tau$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
-Thinking non-interactive Plan flight schedule ($\tau$-bench Airline); consider returning items ($\tau$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
-State-changing tool call (write) non-interactive Update flight booking ($\tau$-bench Airline); return/cancel items ($\tau$-bench Retail); Turn on Wifi (ToolSandbox)
-Providing final information to user non-interactive Summarize updated flights ($\tau$-bench Airline); summarize returned items ($\tau$-bench Retail); summarize phonebook update results (ToolSandbox)
-Table 2: Example action classes in three scenarios from two benchmarks, $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
+Below are the example prompts one can adopt in $\tau^{2}$-bench (Retail) for LLM-judge in interactivity and evidentiality classification.
+Example System Prompt for Interactivity Classification You are classifying customer service agent actions as either: - "interactive": asking questions, requesting clarification, using tools to collect info - "non_interactive": thinking, state-modifying actions, providing final answers Respond as a JSON array of objects with keys: "action_idx" (int), "label" (str), "justification" (str).
+Example System Prompt for Evidentiality Classification You are evaluating whether a customer service agent’s actions are grounded in the conversation context and domain policy.
+For each assistant action, determine: - Are tool call arguments traceable to user-provided information or prior tool results?
+- Are text claims about policies consistent with the provided policy document?
+- Are stated facts (account details, prices, statuses) confirmed by tool results in the conversation?
+Labels: - "evidential": The action is well-grounded in the conversation and policy.
+- "non_evidential": The action contains fabricated information, hallucinated tool arguments, or claims contradicting the policy.
+- "uncertain": Cannot determine grounding with available context.
+Respond as a JSON array of objects with keys: "action_idx" (int), "label" (str), "justification" (str), "issues" (list of str, empty if none).
 Uncertainty estimation.
-In our proposed total uncertainty lower bound Eq.
-6, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty.
-The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or the conformal prediction set size [kumar2023conformal].
-However, the observation uncertainty and the initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world.
+In the proposed total uncertainty lower bound Eq.
+7, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty.
+The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or verbalized confidence methods [tian2023just, xiong2024can, yang2024verbalized].
+However, as mentioned in Section 4.1, all these existing approaches have their own limitations, and one may have to choose an appropriate estimator depending on the situation.
+Meanwhile, observation uncertainty and initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world (See Section 4.2).
 If the system allows multiple trials, one can utilize a sampling-consistency-based approach to estimate the observation uncertainty, likewise the traditional black-box UQ methods.
@@
 Another line of methods can be establishing a world model to approximate the environment transition and estimate observation uncertainty from it [chae2025web, zhou2025dinowm, anonymous2026dreamphase].
+Figure 8 provides results on an auxiliary LLM-based approximation that mimics world modeling philosophy.
 Mutual information estimation.
-The proposed information gating mechanism features the information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$.
+The proposed information gating mechanism features information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and the initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$.
 While the estimation of mutual information is notorious for its difficulty, recent advancements on neural estimators [belghazi2018mutual, cheng2020club, mukherjee2020ccmi, molavipour2020conditional, gritsai2025mist] shed light on this problem by enabling estimation in high-dimensional, unstructured data space.
+Besides, LLM-prompting-based mutual information estimation [robertson2025measure] is also emerging as an attractive line of work, though it has yet to provide a solid theoretical foundation.
 Conditional certainty maximization process.
-As an analogy to the conditional uncertainty reduction process presented in Section 4.2, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix B.
+As an analogy to the conditional uncertainty reduction process presented before, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix C.
 Given that Onicescu’s informational energy brings its correlation coefficient as well, we may be able to implement the exact same style information-gating-based method that has the opposite direction to uncertainty reduction dynamics.
-Appendix E Action Categorization
-Along with the flight booking scenario in Table 1, we extend our action category analysis to two more scenarios, retail and telecommunication tasks, from $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] in Table 2.
-Appendix F Missing Proof
-Lemma 2 (Restatement of Lemma 1).
+E.2 Theoretical Analysis
+Lemma 1 (Extrema of Information Gating).
 Let the lower bound of agent total uncertainty in Eq.
@@
 ∎
-Appendix G Discussion and Future Work
+E.3 Action Categorization
+Action Category Interactivity Example
+Information-gathering interactive Agent uses a read tool (e.g., get_reservation_details or search_flights) to retrieve data or asks the user for missing information, e.g., “Could you provide your reservation number and last name?”.
+Asking clarification or confirmation to user interactive Agent asks the user preferences or confirmation on decision, such as, “Do you want me to proceed with booking FQ8APE?”.
+Thinking non-interactive Agent plans the future action sequence based on the previous trajectory.
+State-changing tool call (writing) non-interactive Agent calls the tools that modify the database (e.g., cancel_reservation, book_reservation, update_reservation_flights) commit to an outcome.
+Providing final information to user non-interactive Agent reports the result of an action, such as “Your reservation has been cancelled; your refund will be processed” or “Your flight has been rebooked to SFO departing at 8 a.m.”.
+Table 3: Example action classes in Airline booking scenario in $\tau^{2}$-bench.
+A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty.
+See Table 4 for extended analyses.
+Action Category Interactivity Scenario-specific Example
+Information-gathering interactive Search flight information ($\tau^{2}$-bench Airline); retrieve order status ($\tau^{2}$-bench Retail); read message box (ToolSandbox)
+Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau^{2}$-bench Airline); instruct or ask the user to choose cancellation ($\tau^{2}$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
+Thinking non-interactive Plan flight schedule ($\tau^{2}$-bench Airline); consider returning items ($\tau^{2}$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
+State-changing tool call (write) non-interactive Update flight booking ($\tau^{2}$-bench Airline); return/cancel items ($\tau^{2}$-bench Retail); Turn on Wifi (ToolSandbox)
+Providing final information to user non-interactive Summarize updated flights ($\tau^{2}$-bench Airline); summarize returned items ($\tau^{2}$-bench Retail); summarize phonebook update results (ToolSandbox)
+Table 4: Example action classes in three scenarios from two benchmarks, $\tau^{2}$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
+An actual example categorization of the action interactivity in an airline assistant task is provided in Table 3.
+Besides, a higher abstraction of action category to two more scenarios, retail in $\tau^{2}$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] is provided in Table 4.
+Appendix F Experiment
+To ground our statements with empirical results, we conducted a small-scale pilot experiment on a real-world agent benchmark in general assistance tasks to evaluate existing UQ techniques on LLM agents.
+F.1 Setup
+Dataset and task.
+$\tau^{2}$-bench [barres2025tau] (the latest version of $\tau$-bench [yao2024tau]) is a representative agent benchmark adopted in frontier LLM evaluation [gemini_team2026gemini31pro], consisting of three different domains – airline, retail, and telecom, where the total number of sample tasks is 50, 114, and 114, respectively.
+By following gemini_team2026gemini31pro, we experimented with retail and telecom domains.
+The tasks in the Retail domain are about helping users cancel or modify pending orders, return or exchange delivered orders, modify user addresses, or provide information.
+The tasks in the telecom domain are about tackling telecommunication issues in mobile devices, organized into roughly three categories: service issues, mobile data issues, and MMS issues.
+LLMs and inference details.
+We experiment with two modern LLMs: GPT-4.1 [openai2025gpt41] and Kimi-K2.5 [kimi2026] through Microsoft Azure [^4] https://azure.microsoft.com/en-us AI Foundry API endpoints.
+Most modern proprietary LLMs do not provide the output log probability in multiple inference engine consistency, and GPT-4.1 was a viable choice for us to adopt here.
+Meanwhile, small-scale open-source models that can be run in local GPUs did not achieve meaningful performance on $\tau^{2}$-bench due to its task complexity, so we stick with a state-of-the-art large-scale open-source model, Kimi-K2.5, through third party platform.
+We set the temperature to 0.0 during inference, which is the default value of $\tau^{2}$-bench.
+Due to the resource constraint, we only run a single trial per task in contrast to the four trials conducted in $\tau^{2}$-bench.
+User simulator LLM was fixed to Kimi-K2.5 for all tasks, and we only varied agent LLM between GPT-4.1 and Kimi-K2.5.
+Implementation of uncertainty estimators.
+We adopt three different uncertainty/confidence estimators: negative log-likelihood (NLL), Entropy, and verbalized confidence.
+The NLL and Entropy are calculated per token and averaged across the entire sequence of actions in the trajectory, whereas the verbalized confidence is calculated per turn and also averaged across the entire trajectory.
+To be specific,
+  • $\hat{U}_{\text{NLL}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{U}_{\text{NLL}}(A_{i}|E_{i-1},O_{i-1})=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}-\log P^{1}_{\pi,\mathcal{T}}(A_{i,j}|E_{i-1},O_{i-1})$
+  • $\hat{U}_{\text{Ent}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{U}_{\text{Ent}}(A_{i}|E_{i-1},O_{i-1})=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}\mathbb{E}_{A_{i,j}}[-\log P^{5}_{\pi,\mathcal{T}}(A_{i,j}|E_{i-1},O_{i-1})]$
+  • $\hat{C}_{\text{verb}}(\mathcal{F}_{\leq T})=\frac{1}{T}\sum_{i=1}^{T}\hat{C}_{\text{verb}}(A_{i}|E_{i-1},O_{i-1},\texttt{vc-prompt})$
+where $L_{i}$ denotes the sequence length of $i$-th turn, $P^{K}$ indicates top-K output token distribution (a chosen token probability if $K=1$), and vc-prompt is a prompt for verbalized confidence we designed as below,
+System Prompt for Verbalized Confidence [tian2023just] Provide the probability (0.0 to 1.0) that your current response/action is correct for the given conversation history and context so far.
+Give ONLY the probability, no other words or explanation.
+Evaluation details.
+Since $\tau^{2}$-bench does not provide turn-level annotation, we conduct trajectory-level evaluation that compares the trajectory-level uncertainty $U(\mathcal{F}_{\leq T})$ with the final task failure 1r1-r (or success $r$ in the case of verbalized confidence).
+For this, we measure the area under the receiver operating characteristic curve (AUROC), Spearman rank correlation coefficient $\rho$, and Kendall rank correlation coefficient $\tau$.
+The average reward (success rate) per model-domain pair is also reported in Table 2.
+F.2 Extended Results
+This subsection provides full analysis results on the observation uncertainty in Figure 8 and uncertainty evolution in Figure 9 and Figure 10.
+In Figure 8, we compare three different observation uncertainties over all user messages across trajectories: (1) the GT user simulator uncertainty (computed from the output probability of user simulator LLM), (2) estimates driven from the output probability of agent LLM by feeding user messages along with all previous context, and (3) estimates driven from the output probability of an auxiliary LLM by feeding user messages along with all previous context, coupled with a prompt as below.
+System Prompt for Observation Distribution Approximator You are an impartial observer monitoring a customer-service conversation between an AI agent and a human user.
+Your task is to predict the next observation that the agent will receive: 1.
+A user message - the customer’s natural-language reply, or 2.
+A tool result - the deterministic output returned by a backend API after the agent issues a tool call.
+Given the full conversation history up to this point, generate the most likely continuation as if you were the user or the tool backend.
+Aim to model the distribution of plausible observations as faithfully as possible: assign high probability to expected responses and low probability to surprising ones.
+Figure 8: Comparison of observation uncertainty estimates on $\tau^{2}$-bench.
+We compare distributions over observation uncertainty estimates, i.e., average NLL and Entropy over all text message tokens from the user.
+Distributions of the ground truth user simulator LLM and that of the agent-approximated one show remarkable deviation, whereas auxiliary LLM prompting-based estimation shows somewhat close distribution with the ground truth.
+We observe remarkable deviation between the GT user simulator’s uncertainties and the estimates from the agent LLM in general.
+However, using an auxiliary LLM as a world model somehow reduces the gap in many cases.
+Cost-effective auxiliary model-based approximation can be worth exploring.
+Next, Figure 9 and 10 compare uncertainty evolutions over turns between success and failure groups of tasks.
+To compare varying-length trajectories across different tasks, we normalize the turn index in X-axis into a zero-to-one scale.
+Results show that existing naive uncertainty aggregation methods can not discern the failure group from the success group to a statistically meaningful degree (the shaded areas in the plots denote a single standard deviation confidence interval), both in the middle of and at the end of the trajectory.
+Figure 9: Uncertainty evolution of GPT-4.1 across the trajectory on $\tau^{2}$-bench.
+We visualize the evolution of action uncertainty/confidence estimates over the normalized turn index, averaged over the entire trajectories from two groups: task success and failure.
+The shade area denotes one standard deviation.
+Figure 10: Uncertainty evolution of Kimi-K2.5 across the trajectory on $\tau^{2}$-bench.
+We visualize the evolution of action uncertainty/confidence estimates over the normalized turn index, averaged over the entire trajectories from two groups: task success and failure.
+The shade area denotes one standard deviation.
+Appendix G Mini-Survey on Agent Benchmark
+We did a small-scale survey on LLM-based agent benchmarks released to the public from February 2023 to February 2026, where we sourced the papers mainly from arxiv.org with non-zero citations, and also referred to ML/NLP conference proceedings.
+We ended up with a non-exhaustive collection of 44 papers, mainly designed for pure LLM agents, while a few of them also support multimodal scenarios of VLM agents.
+We confine the scope of the survey to the agent’s problem-solving capability rather than other desired properties such as safety and trustworthiness [zhang2024agent, levy2024st, andriushchenko2025agentharm].
+We categorize the papers into three classes depending on the evaluation granularity: turn-level, milestone-level, and trajectory-level.
+Table 5 summarizes the turn-level and milestone-level evaluation benchmarks, and Table 6 summarizes the trajectory-level ones.
+The scarcity of fine-grained benchmarks is severe and becomes a real bottleneck to establishing and evaluating the agent UQ framework.
+Title Citation Evaluation
+Agent-as-a-Judge: Evaluate Agents with Agents pmlr-v267-zhuge25a Turn-level
+API-Bank: A Comprehensive Benchmark for Tool-Augmented LLMs li-etal-2023-api Turn-level
+Clembench: Using Game Play to Evaluate Chat-Optimized Language Models as Conversational Agents chalamalasetti2023clembench Turn-level
+Mind2Web: Towards a Generalist Agent for the Web deng2023mind2web Turn-level
+AgentBoard: An Analytical Evaluation Board of Multi-turn LLM Agents chang2024agentboard Milestone-level
+DevBench: A Comprehensive Benchmark for Software Development li2025prompting Milestone-level
+Gaia2: Benchmarking LLM Agents on Dynamic and Asynchronous Environments froger2026gaia2 Milestone-level
+PaperBench: Evaluating AI’s Ability to Replicate AI Research pmlr-v267-starace25a Milestone-level
+ScienceAgentBench: Toward Rigorous Assessment of Language Agents for Data-Driven Scientific Discovery chen2025scienceagentbench Milestone-level
+ST-WebAgentBench: A Benchmark for Evaluating Safety and Trustworthiness in Web Agents levy2024st Milestone-level
+The Berkeley Function Calling Leaderboard (BFCL): From Tool Use to Agentic Evaluation of Large Language Models patil2025berkeley Milestone-level
+TheAgentCompany: Benchmarking LLM Agents on Consequential Real World Tasks xu2025theagentcompany Milestone-level
+ToolSandBox: A Stateful, Conversational, Interactive Evaluation Benchmark for LLM Tool Use Capabilities lu2025toolsandbox Milestone-level
+WebCanvas: Benchmarking Web Agents in Online Environments pan2024webcanvas Milestone-level
+Table 5: Agent benchmark research with turn-level and milestone-level evaluation protocols.
+Title Citation
+AgentBench: Evaluating LLMs as Agents liu2024agentbench
+AgentGym: Evolving Large Language Model-based Agents across Diverse Environments xi-etal-2025-agentgym
+AppWorld: A Controllable World of Apps and People for Benchmarking Interactive Coding Agents trivedi-etal-2024-appworld
+AssistantBench: Can Web Agents Solve Realistic and Time-Consuming Tasks? yoran2024assistantbench
+BrowseComp: A Simple Yet Challenging Benchmark for Browsing Agents wei2025browsecomp
+CORE-Bench: Fostering the Credibility of Published Research Through a Computational Reproducibility Agent Benchmark siegel2024corebench
+Finance Agent Benchmark: Benchmarking LLMs on Real-world Financial Research Tasks bigeard2025finance
+GAIA: A Benchmark for General AI Assistants mialon2023gaia
+GameBench: Evaluating Strategic Reasoning Abilities of LLM Agents costarelli2024gamebench
+InterCode: Standardizing and Benchmarking Interactive Coding with Execution Feedback yang2023intercode
+MCPAgentBench: A Real-world Task Benchmark for Evaluating LLM Agent MCP Tool Use liu2025mcpagentbench
+MCP-Atlas: A Large-Scale Benchmark for Tool-Use Competency with Real MCP Servers bandi2026mcpatlas
+MedAgentBench: A Realistic Virtual EHR Environment to Benchmark Medical LLM Agents jiang2025medagentbench
+MINT: Evaluating LLMs in Multi-turn Interaction with Tools and Language Feedback wang2024mint
+MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation pmlr-v235-huang24y
+OSWorld: Benchmarking Multimodal Agents for Open-Ended Tasks in Real Computer Environments xie2024osworld
+SafeAgentBench: A Benchmark for Safe Task Planning of Embodied LLM Agents yin2024safeagentbench
+SmartPlay: A Benchmark for LLMs as Intelligent Agents wu2024smartplay
+SOTOPIA: Interactive Evaluation for Social Intelligence in Language Agents zhou2024sotopia
+SWE-bench: Can Language Models Resolve Real-World GitHub Issues? jimenez2024swebench
+Terminal-Bench: Benchmarking Agents on Hard, Realistic Tasks in Command Line Interfaces merrill2026terminal
+The Tool Decathlon: Benchmarking Language Agents for Diverse, Realistic, and Long-Horizon Task Execution li2025tooldecathlon
+ToolLLM: Facilitating Large Language Models to Master 16000+ Real-world APIs qin2024toolllm
+TravelPlanner: A Benchmark for Real-World Planning with Language Agents xie2024travelplanner
+Vending-Bench: A Benchmark for Long-Term Coherence of Autonomous Agents backlund2025vending
+WebArena: A Realistic Web Environment for Building Autonomous Agents zhou2024webarena
+WebShop: Towards Scalable Real-World Web Interaction with Grounded Language Agents yao2022webshop
+WorkArena: How Capable Are Web Agents at Solving Common Knowledge Work Tasks? drouin2024workarena
+$\tau$-bench: A Benchmark for Tool-Agent-User Interaction in Real-World Domains yao2024tau
+$\tau^{2}$-Bench: Evaluating Conversational Agents in a Dual-Control Environment barres2025tau
+Table 6: Agent benchmark research with a trajectory-level evaluation protocol.
+Appendix H Discussion and Future Work
 Does solely establishing the agent UQ framework actually help us to reliably infer on the agent’s performance on a task?